Bernie lives in Boston. Tina lives 1175 miles away in Florida. At 9:00 am on December 24th, Bernie heads south at a speed of 51 mph. At 11:00 am on December 24th, Tina heads north at a speed of 65 mph. What time will Bernie and Tina pass each other? What is the formula?
Found 2 solutions by TimothyLamb, richwmiller:
Answer by TimothyLamb(4379)

---
Bernie's head start:
s = d / t
d = s * t
t = 11 - 9 = 2 hours
d = 51*2
d = 102 miles
---
distance between them after Bernie's head start:
1175 - 102 = 1073 miles
---
relative speed of Bernie and Tina:
s = 51 + 65
s = 116 mph
---
time to their meeting, relative to Tina's departure:
t = d / s
t = 1073/116
t = 9.25 hours
---
clock time:
11 + 9.25 = 20.25
---
answer:
they will pass each other at 8:15 PM
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by richwmiller(17219)

Bernie traveled 2*51=102 miles before Tina left Florida.
1175-102=1073
51t+65t=1175-2*51
106t=1073
t=10.122641509433962264150943396226
t=10 hours 7 minutes 21 seconds
9:07:21 pm